# Prospects for Observing High-redshift Radio-loud Quasars in the SKA Era: Paving the Way for 21-cm Forest Observations

### v1

@@ 2.3 Accretion Rate and Optical Luminosity
where $L_{\rm Edd}$ is the Eddington luminosity. As the discussion above, we adopt $\lambda_{\rm Edd}=1$ for most of our calculations. As an additional discussion, we also calculated the case of $\lambda_{\rm Edd}=1.2$.
For quasar bolometric luminosity corrections, we use the results of Runnoe et al. (2012) (their Equations (11) and (13)),
@@
We use the average spectral energy distribution from Shang et al. (2011) to calculate the luminosity relationship between different optical bands.
@@ 2.4 Duty Cycle of Quasar Phase
Through the halo mass function and the halo-SMBH mass relation, we have obtained the abundance of black holes. However, not all black holes are in the quasar phase at the same time. We need to estimate the proportion of black holes that are in the quasar phase, which is the duty cycle. We estimate the duty cycle as
@@
The average lifetime $t_{\rm q}$ and average accretion rate $\lambda_{\rm Edd}$ of quasars are strongly correlated parameters. We obtained the best fit to the observed luminosity function data by varying the average lifetime $t_{\rm q}$ of the quasars, using a fixed average Eddington ratio $\lambda_{\rm Edd}=1$ for early quasars. We employed two sets of high-redshift observational data, obtained from observations of $z\sim 6$ (Matsuoka et al., 2018) and $z\sim 6.8$ (Matsuoka et al., 2023). Our best-fit results are shown in Figure 2, and we take the logarithmic mean value $10^{5.275}$ yr in subsequent calculations. Note that both sets of data, especially the $z\sim 6.8$ data, are affected by incomplete observation coverage at the faint end, so we did not use the first two data points at the faint end in the fitting process. As a possible scenario, we also calculated the fit with an average Eddington ratio of $\lambda_{\rm Edd}=1.2$, which is shown in Figure 3.
@@
Figure 3: Similar to Figure 2. The average Eddington ratio of quasars is fixed at 1.2.
@@ 2.5 Radio-Loud Fraction
where $M_{2500}$ represents the optical magnitude at 2500Å. $b_{0}$, $b_{z}$ and $b_{M}$ are free parameters, with best-fit values of $-0.132$, $-2.052$, and $-0.183$, respectively. Since the recent study on the RLF of $z\geq 6$ quasars does not rule out the possibility of this form of evolution (Keller et al., 2024), we also calculated its impact on the results (see Section 3.5).
@@ 2.6 Radio-Loudness Distribution
We used the radio loudness data of 225 radio-loud quasars with redshift from 0 to 5 from Bariuan et al. (2022) to perform a Gaussian fit of the radio-loudness distribution. The definition of radio-loudness is the logarithm of the ratio of the rest frame luminosity between 5 GHz and 4400Å ($R={\rm log_{10}}\left(L_{\rm 5GHz}/L_{4400}\right)$). We estimate the radio-loudness distribution of the radio loud quasars to be
@@
where $R$ is radio-loudness, and $\bar{R}$ and $\sigma_{0}$ are free parameters. Figure 4 shows the result of our radio-loudness distribution fitting, with the best fitting result being $\bar{R}=2.67$ and $\sigma_{0}=0.55$. We also performed fits for the high-redshift and low-redshift data separately and did not find any significant evidence of redshift evolution. Therefore, we believe that applying this distribution form to higher redshifts should be safe.
Figure 4: The fitting results of the radio-loudness distribution. The radio-loudness data used are from Bariuan et al. (2022). The red line represents our best fit to the data, the orange line represents the radio-loudness distribution H04 used.
@@ 2.7 Radio Spectral Index
where $F_{\rm radio}$ is the radio flux and $\alpha$ is the radio spectral index. The radio spectral index affects two calculation processes: one involves converting from the observational frame to the rest frame, known as K-correction, and the other is the transformation between the 5 GHz definition of radio-loudness and the SKA-LOW observing band. Based on the findings of Gloudemans et al. (2021), who derived a median spectral index of $-0.29$ through the stacking of 93 quasars using Low Frequency Array Two Metre Sky survey (LoTSS-DR2) and Faint Images of the Radio Sky at Twenty Centimeters (FIRST) data, we adopt $\alpha=-0.29$. Note that previous high-redshift studies based on a small number of observational data used a steeper spectral index of $\alpha=-0.75$. Our spectral index value is supported by a larger sample (93 $z>5$ quasars), which should be more reliable. Due to our limited understanding of high-redshift quasar radio spectra, we also calculated the impact of $\alpha=-0.75$ on the results as a possible scenario (see Section 3.4).
@@ 2.8 Radio-Loud Quasar Abundance
From the above definition, we can get the radio-loud quasar abundance with flux greater than a certain flux threshold $F_{\rm radio}$ as
$\frac{{\rm d}N}{{\rm d}z{\rm d}\Omega}(F_{\rm radio},z)=\frac{{\rm d}V}{{\rm d}z{\rm d}\Omega}D_{\rm q}$
×$\times\quad\int_{0}^{\infty}{\rm d}M_{\rm halo}\left(\int_{R_{0}(M_{\rm halo},F_{\rm radio})}^{\infty}{\rm d}R\ N(R)\right)\frac{{\rm d}n}{{\rm d}M_{\rm halo}},$ (10)
where ${\rm d}V/{\rm d}z{\rm d}\Omega$ is the cosmological volume element, $R_{0}$ is the threshold of radio loudness calculated from the given radio flux density threshold $F_{\rm radio}$ and halo mass $M_{\rm halo}$, and ${\rm d}n/{\rm d}M_{\rm halo}$ is the halo abundance calculated from Equation (1).
@@ 3.1 Quasar Abundance Predicted by The Model
Figure 5 shows the distribution of quasar abundance above a given radio flux threshold as a function of redshift, obtained according to Equation (10). As discussed in Section 2, the standard parameter settings we used are as follow: average Eddington ratio $\lambda_{\rm Edd}=1$, average quasar lifetime $t_{\rm q}=10^{5.275}$ years, radio spectral index $\alpha=-0.29$, and radio-loudness fraction RLF = 10%.
Figure 5: Predicted distribution of radio-loud quasar abundance above different flux thresholds as a function of redshift, with the observation band at 150 MHz. The characteristic abundance ${\Phi}^{*}$ corresponds to an abundance where the expected number of observations over the entire sky ($\sim 41253\ {\rm deg}^{2}$) is 1 (${\Phi}^{*}\simeq 1/41253\ {\rm deg}^{-2}$). Sources with an abundance below this value are almost non-existent.
@@ 3.2 Comparison with LOFAR Observations
We attempt to compare the calibrated model with existing high-redshift radio observations. Specifically, we use data from the EALIS-N1 field in the LoSST Deep Fields DR1 (Sabater et al., 2021; Best et al., 2023), which has the lowest average noise level in the three deep fields. We select radio-excess AGNs with redshifts between 5.6 and 6.4 from this deep field observation data and construct the radio luminosity function using the standard $1/V_{\rm max}$ technique (Schmidt, 1968; Condon, 1989). The detailed calculation process and incompleteness information can be found in Section 3.1 of Kondapally et al. (2022). For the model, we convolve the optical luminosity function with the radio-loudness distribution to calculate the radio luminosity function. Figure 6 shows the comparison of the radio luminosity function predicted by the model with the LOFAR observations. At $z\sim 6$, the model-predicted radio luminosity function matches well with the LOFAR observations. However, we also notice that the observed radio luminosity function is slightly higher than the model prediction (within the error range), and this discrepancy is more pronounced in the Lockman Hole field (exceeding the $1\sigma$ error range). Since our model is calibrated using the observed optical luminosity function, this discrepancy may indicate that some high accretion rate AGNs during the reionization epoch are optically obscured, which is consistent with some recent high-redshift researches (Endsley et al., 2023; Lambrides et al., 2024; Banados et al., 2024).
Figure 6: Comparison of the radio luminosity function with observational data. The blue line represents the radio luminosity function of radio-loud quasars at $z=6$ calculated by the model, while the red data points are the observational results from the EALIS-N1 field in the LoSST Deep Fields DR1, with error bars representing the $1\sigma$ error.
@@ 3.3 SKA-LOW Observation Forecast
For SKA-LOW radio-loud quasar observations, we considered two observation modes: the first is a survey to search for quasars. Since the goal of this survey is to find visible radio-loud quasars in the sky, a high frequency resolution is not required, so we refer to it as a continuum survey. The relevant parameters for the continuum survey are summarized in Table 1. The second mode is a survey to find radio-loud quasars suitable for 21-cm forest studies. To resolve individual 21-cm lines, this survey requires higher frequency resolution, and thus we call it the 21-cm forest survey. The relevant parameters for the 21-cm forest survey are summarized in Table 2. Note that our 21-cm forest survey forecast shows the total number of sources within a survey area of 10313 ${\rm deg}^{2}$ that can resolve individual 21-cm lines with 100 h of integration time. Conducting a 21-cm forest blind survey with this setup would require a substantial amount of observation time ($\sim$ 9 years). From an observational perspective, we should first conduct the continuum blind survey to determine the positions of visible sources and then perform a high frequency-resolution follow-up observations on those bright sources. How to filter high-redshift candidates from the blind survey results is an important issue for future observations, which is not discussed in this paper.
Figure 7: Forecast of the SKA-LOW observations. Dash lines represent continuum survey, and solid lines represent 21-cm forest survey. The parameters for continuum survey and 21-cm forest survey are shown in Table 1 and Table 2, respectively.
$\delta t=\frac{T_{\rm tot}}{S_{\rm tot}/{\mathit{\Omega}}},$ (12)
where ${\mathit{\Omega}}$ is the field of view, given by ${\mathit{\Omega}}\simeq\left(1.2\lambda/D_{\rm ant}\right)^{2}$. We estimate ${\mathit{\Omega}}$ using the wavelength corresponding to 150 MHz. We assume that the signal-to-noise ratio is greater than 5 to detect quasars. For continuum survey, this means ${\rm Flux}>5\times\delta{S}^{\rm{N}}$. For 21-cm forest survey, we assume a optical depth of $\tau=0.1$ on the basis of a signal-to-noise ratio of 5, which means ${\rm Flux}\times\left(1-e^{-\tau}\right)>5\times\delta{S}^{\rm{N}}$.
@@
Survey Time $T_{\rm tot}$ $365\times 6$ hours
@@
Integration Time $\delta t$ 100 hours

### v2

@@ 2.3 Accretion Rate and Optical Luminosity
where $L_{\rm Edd}$ is the Eddington luminosity. As the discussion above, we adopt $\lambda_{\rm Edd}=1$ for most of our calculations. We will also discuss the case with $\lambda_{\rm Edd}=1.2$.
For the quasar bolometric luminosity corrections, we use the results from Runnoe et al. (2012) (their Equations (11) and (13)):
@@
We adopt the average spectral energy distribution (SED) from Shang et al. (2011) to calculate the luminosity relationship between different optical bands:
$\lambda F_{\lambda}\left(1450\AA\right)\simeq 1.86\ \lambda F_{\lambda}\left(5100\AA\right),$ (8)
$\lambda F_{\lambda}\left(2500\AA\right)\simeq 1.76\ \lambda F_{\lambda}\left(5100\AA\right),$ (9)
$\lambda F_{\lambda}\left(4400\AA\right)\simeq 1.11\ \lambda F_{\lambda}\left(5100\AA\right),$ (10)
where $\lambda F_{\lambda}(\lambda)$ is the rest-frame spectral flux at wavelength $\lambda$.
@@ 2.4 Duty Cycle of Quasar Phase
Through the halo mass function and the halo–SMBH mass relation, we have obtained the black hole mass function. However, not all black holes are in the quasar phase at the same time. We need to estimate the proportion of black holes that are in the quasar phase, which is the duty cycle. We estimate the duty cycle as
@@
By combining the black hole mass function with Equations (6)–(8) and (11), we can calculate the quasar luminosity function (QLF, the comoving number density of quasars per unit of magnitude) as
$\Phi\left(M_{1450},z\right)=\frac{{\rm d}n_{\rm q}}{{\rm d}M_{1450}}=D_{\rm q}\frac{{\rm d}n_{\rm BH}}{{\rm d}M_{\rm BH}}\frac{{\rm d}M_{\rm BH}}{{\rm d}M_{1450}},$ (12)
where $n_{\rm q}$ is the comoving number density of quasars, ${\rm d}n_{\rm BH}/{\rm d}M_{\rm BH}$ is the black hole mass function, and $M_{1450}$ is the absolute magnitude at $1450\ \AA$,
$M_{1450}=-2.5{\rm log}\left(\frac{L_{\nu}\left(\nu(1450\AA)\right)}{4\pi d^{2}3631{\rm\ Jy}}\right),$ (13)
where $d=10\ {\rm pc}$ is expressed in cm.
In the calculation of QLF, $\lambda_{\rm Edd}$ and $t_{\rm q}$ are two parameters that are not strictly constrained by observations. Since both jointly determine the shape of QLF, they are strongly correlated parameters when fitting QLF to observations. Based on the discussion of the accretion rate in Section 2.3, we obtained the best fit to the observed luminosity function data by varying the average lifetime $t_{\rm q}$ of the quasars until $\chi^{2}$ is minimized, using a fixed average Eddington ratio $\lambda_{\rm Edd}=1$ for early quasars. We employed two sets of high-redshift observational data, obtained from observations of $z\sim 6$ (Matsuoka et al., 2018) and $z\sim 6.8$ (Matsuoka et al., 2023). Our best-fit results are shown in Figure 2. The reduced $\chi^{2}$ obtained from the fits at $z=6$ and $z=6.8$ are 1.49 and 0.79, respectively. We take the logarithmic mean value $10^{5.5}$ yr in subsequent calculations. Note that both sets of data, especially the $z\sim 6.8$ data, are affected by incomplete observation coverage at the faint end, so we did not use the data points at the faint end with $M_{1450}>-24$ in the fitting process.
Figure 3: Similar to Figure 2, but the average Eddington ratio of quasars is fixed at 1.2.
As a possible scenario, we also calculated the fit with an average Eddington ratio of $\lambda_{\rm Edd}=1.2$, which is shown in Figure 3. The reduced $\chi^{2}$ obtained from the fits at $z=6$ and $z=6.8$ are 1.26 and 0.68, respectively. From the results, it seems that the fits with a higher $\lambda_{\rm Edd}$ is better, but we must point out that this result could still be influenced by the incomplete observations of high-redshift QLF. The incompleteness effect can be divided into two parts. First, the sensitivity limitation, where there is incompleteness in the data at the faint end of QLF, and we mitigate this effect by discarding the data points with $M_{1450}>-24$. Second, the obscuration by the AGN dust torus may affect the completeness of optical observations, meaning that some AGNs cannot be effectively detected in the optical band.
@@ 2.5 Radio-loud Fraction
where $M_{2500}$ represents the optical magnitude at $2500\ \AA$. $b_{0}$, $b_{z}$ and $b_{M}$ are free parameters, with best-fit values of $-0.132$, $-2.052$, and $-0.183$, respectively. Since the recent study on the RLF of $z\geq 6$ quasars does not rule out the possibility of this form of evolution (Keller et al., 2024), we also calculated its impact on the results (see Section 3.6).
Figure 4: The fitting results of the radio-loudness distribution. The radio-loudness data used are from Bariuan et al. (2022). The red line represents our best fit to the data, and the orange line represents the radio-loudness distribution used in H04.
@@ 2.6 Radio-loudness Distribution
The radio-loudness distribution of AGNs shows a bimodal distribution in observations (e.g. Strittmatter et al., 1980; Kellermann et al., 1989; Ivezić et al., 2002; White et al., 2007), suggesting the presence of two populations with different radio properties. The distribution can be well fitted using two Gaussian components (e.g. Baloković et al., 2012; Xiao et al., 2022). In this study, we focus on the radio-loud population. We used the radio-loudness data of 225 radio-loud quasars with redshift from 0 to 5 from Bariuan et al. (2022) to perform a Gaussian fit of the radio-loudness distribution. The definition of radio-loudness is the logarithm of the ratio of the rest-frame luminosity between 5 GHz and $4400\ \AA$  ($R={\rm log_{10}}\left(L_{\rm 5GHz}/L_{4400}\right)$). We estimate the radio-loudness distribution of the radio-loud quasars to be
@@
where $R$ is radio-loudness, and $\bar{R}$ and $\sigma_{0}$ are free parameters. Figure 4 shows the result of our radio-loudness distribution fitting, with the best-fit result being $\bar{R}=2.67$ and $\sigma_{0}=0.55$. We also divided the data into four redshift bins for individual fitting [^2] The four redshift bins are $[0,1.5]$, $(1.5,2.5]$, $(2.5,3.5]$, and $[3.5,5]$., with the results presented in Figure 5. As indicated, no significant redshift evolution is observed across the four bins. Separate Kolmogorov-Smirnov test (K-S test) was performed for every two redshift bins. All groups, except for the redshift range $(2.5,3.5]$, passed the K-S test. These results suggest that radio-loudness $R$ either remains constant or exhibits only weak evolution with redshift $z$. Therefore, we believe that applying this distribution form to higher redshifts should be safe.
Figure 5: The radio-loudness fitting results for the four redshift bins. Green data points represent the means obtained from Gaussian fitting, with error bars indicating the $\sigma_{0}$ values. The red dashed line and shaded area denote the best fit, as depicted in Figure 4.
@@ 2.7 Radio Spectral Index
where $F_{\rm radio}$ is the radio flux and $\alpha$ is the radio spectral index. The radio spectral index affects two calculation processes: one involves converting from the observed frame to the rest frame, known as K-correction, and the other is the transformation between the 5 GHz definition of radio-loudness and the SKA-LOW observing band. Based on the findings of Gloudemans et al. (2021), who derived a median spectral index of $-0.29$ through the stacking of 93 quasars using the Low Frequency Array Two Metre Sky survey (LoTSS-DR2) and the Faint Images of the Radio Sky at Twenty Centimeters (FIRST) data, we adopt $\alpha=-0.29$. We also note that previous studies of high-redshift radio quasars used a steeper spectral index with $\alpha=-0.75$ (Wang et al., 2007; Bañados et al., 2015; Liu et al., 2021), which is in agreement with very long baseline interferometry observations (Frey et al., 2005; Momjian et al., 2008; Frey et al., 2011; Momjian et al., 2018). Due to our limited understanding of high-redshift quasar radio spectra, we also calculated the impact of $\alpha=-0.75$ on the results as a possible scenario (see Section 3.5).
Figure 6: Predicted abundance of radio-loud quasars as a function of redshift, with different flux density thresholds at 150 MHz as indicated in the legend. The red line shows the characteristic abundance ${\Phi}^{*}$, which corresponds to an abundance where the expected number of observations over the entire sky ($\sim 41253\ {\rm deg}^{2}$) is 1 (${\Phi}^{*}\simeq 1/41253\ {\rm deg}^{-2}$). Sources with an abundance below this value are almost non-existent.
@@ 2.8 Radio-loud Quasar Abundance
Combining the above prescriptions, we can get the radio-loud quasar abundance with a radio flux density greater than a certain flux density threshold $F_{\rm threshold}$ as
$\frac{{\rm d}N}{{\rm d}z{\rm d}\Omega}(F_{\rm threshold},z)=\frac{{\rm d}V}{{\rm d}z{\rm d}\Omega}D_{\rm q}$
×$\times\quad\int_{0}^{\infty}{\rm d}M_{\rm halo}\left(\int_{R_{0}(M_{\rm halo},F_{\rm threshold})}^{\infty}{\rm d}R\ N(R)\right)\frac{{\rm d}n}{{\rm d}M_{\rm halo}},$ (17)
where ${\rm d}V/{\rm d}z{\rm d}\Omega$ is the cosmological volume element, $R_{0}$ is the threshold of radio-loudness calculated from the given radio flux density threshold $F_{\rm threshold}$ and halo mass $M_{\rm halo}$, duty cycle $D_{\rm q}$ can be found in Equation (11), radio-loudness distribution $N(R)$ can be found in Equation (15), and ${\rm d}n/{\rm d}M_{\rm halo}$ is the halo abundance calculated from Equations (1) and (2).
@@ 3.1 Quasar Abundance Predicted by The Model
Figure 6 shows the distribution of quasar abundance above a given radio flux density threshold as a function of redshift, obtained according to Equation (17). As discussed in Section 2, the standard parameter settings we used are as follow: average Eddington ratio $\lambda_{\rm Edd}=1$, average quasar lifetime $t_{\rm q}=10^{5.5}$ years, radio spectral index $\alpha=-0.29$, and radio-loud fraction RLF = 10%.
Figure 7: Comparison of the model-predicted radio luminosity function with observational data. The blue line represents the radio luminosity function of radio-loud quasars at $z=6$ calculated by our model, while the red data points are the observational results from the ELAIS-N1 field in the LoTSS Deep Fields DR1, with error bars representing the $1\sigma$ error.
@@ 3.2 Comparison with LOFAR Observations
[…]
Figure 8: The number density of radio-loud quasars as expected in SKA-LOW observations. The dashed line represents the number density for a continuum survey, and the solid line represents the available bright sources for 21-cm forest observations. The sensitivity of the 21-cm forest survey shown in the figure corresponds to the flux density limit at $z=9$, while at $z=6$, the sensitivity is $\sim 1.3$ mJy. The parameters for the continuum survey and the 21-cm forest survey are shown in Table 1 and Table 2, respectively.
Figure 9: The integrated number of radio-loud quasars at various redshifts as expected in SKA-LOW observations. The darker blue and orange histograms are the results of the fiducial model with $\alpha=-0.29$, and the lighter blue and gray histograms represent the case where the radio spectral index $\alpha=-0.75$. The redshift bin width is $0.5$. The sensitivity of the 21-cm forest survey shown in the figure corresponds to the flux density limit at $z=9$, while at $z=6$, the sensitivity is $\sim 1.3$ mJy.
@@ 3.3 SKA-LOW Observation Forecast
For SKA-LOW radio-loud quasar observations, we considered two observation modes: the first is a survey to search for quasars. Since the goal of this survey is to find visible radio-loud quasars in the sky, a high frequency resolution is not required, so we refer to it as a continuum survey. The relevant parameters for the continuum survey are summarized in Table 1. The second mode is a survey to find radio-loud quasars suitable for 21-cm forest studies. To resolve individual 21-cm lines, this survey requires higher frequency resolution, and thus we call it the 21-cm forest survey. The relevant parameters for the 21-cm forest survey are summarized in Table 2. Note that our 21-cm forest survey forecast shows the total number of sources within a survey area of 10313 ${\rm deg}^{2}$ that can resolve individual 21-cm lines with 100 h of integration time. Conducting a 21-cm forest blind survey with this setup would require a substantial amount of observation time ($\sim$ 9 years). From an observational perspective, we should first conduct the continuum blind survey to determine the positions of visible sources and then perform a high frequency-resolution follow-up observations on those bright sources. How to filter high-redshift candidates from the blind survey is an important issue for future observations, we have performed the relevant discussions in Section 3.4.
@@
Survey Time $T_{\rm tot}$ $365\times 6$ hour
@@
Integration Time $\delta t$ 100 hour
@@
$\delta t=\frac{T_{\rm tot}}{S_{\rm tot}/{\mathit{\Omega}}},$ (19)
where ${\mathit{\Omega}}$ is the field of view, given by ${\mathit{\Omega}}\simeq\left(1.2\lambda/D_{\rm ant}\right)^{2}$, in which $D_{\rm ant}$ is the diameter of a station. We estimate ${\mathit{\Omega}}$ using the wavelength corresponding to 150 MHz. We require the signal-to-noise ratio to be higher than 5 to detect quasars in a continuum survey, or to detect an absorption line for a follow-up 21-cm observation. For continuum surveys, this means ${\rm flux\ density}>5\times\delta{S}^{\rm{N}}$. For 21-cm forest surveys, we assume an optical depth of $\tau=0.1$, as typical for absorption lines from minihalos or from the IGM with moderate heating during reionization (e.g. Furlanetto 2006; Xu et al. 2011; Ciardi et al. 2013, which means ${\rm flux\ density}>5\times\delta{S}^{\rm{N}}/\left(1-e^{-\tau}\right)$.
Figure 10: The infrared magnitude limit of quasars as required to have a radio flux density exceeding the threshold ($\sim 3\rm mJy$ at $z=9$) for the 21-cm forest survey. The blue and gray lines represent quasars that exceed the average radio-loudness $\bar{R}$ by $1\sigma_{0}$ and $2\sigma_{0}$, respectively. The red and green lines correspond to the $5\sigma$ sensitivity of the Euclid and Roman Space Telescope surveys. The emission line structure is generated by the SED we used. Even quasars with extreme radio-loudness can be detected by future infrared surveys.
3.4 Infrared Limit Estimation
For the actual observation of high-redshift quasars, obtaining redshift information is crucial. Currently, the search for quasars primarily occurs in the optical and near-infrared bands. Candidate objects are selected through color-color diagram cuts (e.g. Jiang et al., 2016; Bañados et al., 2016; Yang et al., 2023; Lin et al., 2024), and redshift values are determined via photometric observations or spectral line identification. For the approximately 20 radio-loud quasars with $z\sim 9$ caculated in Section 3.3, the ideal detection method involves infrared-band photometric surveys, which provide redshift data through the identification of the Lyman-$\rm\alpha$ break. While the JWST offers high sensitivity, it is not capable of conducting wide-field survey. In contrast, current and future space-based infrared telescopes, such as Euclid (Euclid Collaboration et al., 2022) and the Nancy Grace Roman Space Telescope (Roman Space Telescope, Spergel et al., 2015), hold significant potential for such surveys.
Based on the radio flux density threshold for the 21-cm forest survey at $z=9$ obtained in Section 3.3, we calculated the corresponding infrared AB magnitude threshold by considering radio-loudness and quasar SED. The results, along with the survey sensitivities of Euclid and Roman Space Telescope, are presented in Figure 10. Note that the emission line structure is based on the quasar SED we used, and we also fitted the shape of the continuum. As Figure 10 shows, Euclid can detect quasars with radio-loudness close to $\bar{R}+1\sigma_{0}$ over a large survey area ($\sim 14000\ \rm{deg}^{2}$). In contrast, although the Roman Space Telescope covers a smaller area ($\sim 2200\ \rm{deg}^{2}$), it can detect a small number of quasars with extreme radio-loudness ($R\gtrsim\bar{R}+2\sigma_{0}$). Therefore, we believe that in the SKA era, combined with these infrared space telescopes, we can effectively search for high-redshift radio-loud quasars.
Figure 11: Influence of different radio spectral index settings on the abundance of radio-loud quasars. Different colors of lines represent different spectral indices, while different line styles represent different radio flux density thresholds (at 150 MHz). Sources with an abundance below the characteristic abundance marked by the red line are almost non-existent.
Figure 12: Influence of different radio spectral index settings on the forecast results of SKA-LOW observations. The observation frequency band of SKA-LOW is used for calculations here. Specifically, for the continuum surveys, we used 150 MHz, while for the 21-cm forest survey, we used the frequency of the redshifted 21-cm line.